The total number of neutrons present in 54 mL H20 (l) are?:
1
Expert's answer
2020-11-14T13:51:31-0500

Molecular weight of H2​O=18g

No. of neutrons in 1 mole i.e. 18g of H2

​O=8NA

​  ( from O atom, as H has no neutron)


Given moles of H2

​O=18/54 =3


∴Total no. of neutrons present=3×8NA

​=24NA